A loan will be paid by means of payment of 250 each,every 6 months for 10 years.An interest rate of 5% per year compounded every 6 months will be applicable the present value of the loan is


1
Expert's answer
2022-03-07T21:51:01-0500

Solution


Consider,


"A" – The payment per period.


"P" – The principal amount (loan taken).


"r" – The interest rate per year,


"n" – The number of times, the amount is compounded per annum, and


"t" – The number of years,


we can write the formula, as 


"{\color{Blue} P=A\times\frac{(1+\frac{r}{n})^{nt}-1}{(\frac{r}{n})(1+\frac{r}{n})^{nt}}}"


Replacing the values,


"A = 250" ,        "r = 5\%" ,            "n = 2" (compounded every six months), "t = 10" years


"{\color{Red} P=(250)\times\frac{(1+\frac{0.05}{2})^{(2)(10)}-1}{(\frac{0.05}{2})(1+\frac{0.05}{2})^{(2)(10)}}}"


"P=3897.290572\"


Hence the loan amount to be paid will be


"P=3897.290572\"